When the price of Apricot was Rs. 100 per kg, 200kgs of Apricot were

demanded in Delhi market per day. Now the Apricot price has increased to

Rs. 110 per kg as a result of which the quantity demanded reduced to 160

kg per day. Calculate the price elasticity of demand by using the mid - point

method.


1
Expert's answer
2021-03-01T07:25:46-0500

Solution to question #166461


Price Elasticity of demand using mid-point formula.


"Elasticity=\frac{percentage change in quantity}{percentage change in price}"


"Percentage change in quantity =\frac{Q2-Q1}{\frac{Q2+Q1}{2}}\times100%"


"Percentage change in price =\frac{P2-P1}{\frac{P2+P1}{2}}\times100"


"Percentage change in quantity =\frac{160-200}{\frac{160+200}{2}}\times 100=-22.22"


"Percentage change in price=\frac{110-100}{\frac{110+100}{2}}\times 100=9.524"


"Elasticity=\frac{-22.22}{9.524}=-2.333"